# Scene-Text Grounding for Text-Based Video Question Answering

### v1

@@ VI Conclusion
We explore a novel task called Grounded TextVideoQA. The task is crucial for developing trustworthy VideoQA systems involving scene texts. The task also allows for better evaluation of models by enabling direct assessment of grounded scene text regions instead of just textual answers. To solve the task, we propose T2S-QA, a visually grounded TextVideoQA model featuring a disentangled temporal-to-spatial contrastive grounding mechanism for scene text grounding and grounded TextVQA. Additionally, to facilitate evaluation, we construct the ViTXT-GQA benchmark, which enables direct evaluation of spatio-temporal scene-text grounding. Comparative results with existing models and ablated variants demonstrate the superiority and effectiveness of our approach in both QA and answer grounding. Moreover, we conduct an in-depth analysis of the model’s specific behavior concerning grounding and QA accuracy. The results indicate that the primary reasons for poor TextVideoQA performance are the sub-optimal OCR system and ineffective evaluation metrics. We hope our dataset, baselines, and analyses will contribute to the advancements in faithfully answering scene-text questions in videos.

### v2

@@ VI Conclusion
We explore a novel task called Grounded TextVideoQA. The task is crucial for developing trustworthy VideoQA systems involving scene texts. The task also allows for better evaluation of models by enabling direct assessment of grounded scene text regions instead of just textual answers. To solve the task, we propose T2S-QA, a visually grounded TextVideoQA model featuring a disentangled temporal-to-spatial contrastive grounding mechanism for scene text grounding and grounded TextVQA. Additionally, we construct the ViTXT-GQA benchmark to facilitate evaluation, which enables direct evaluation of spatio-temporal scene-text grounding. Comparative results with existing models and ablated variants demonstrate the superiority and effectiveness of our approach in both QA and answer grounding. Moreover, we conduct an in-depth analysis of the model’s specific behavior concerning grounding and QA accuracy. The results indicate that the primary reasons for poor TextVideoQA performance are the sub-optimal OCR system and ineffective evaluation metrics. We hope our dataset, baselines, and analyses will contribute to the advancements in faithfully answering scene-text questions in videos.